# PaDeLLM-NER: Parallel Decoding in Large Language Models for Named Entity Recognition

### v4

@@ 4.2 Main Results
AutoReg CoNLL03 ACE05 GENIA Weibo MSRA Onto4 Resume Youku Ecom Mean
@@
English Dataset Chinese Dataset
AutoReg CoNLL03 ACE05 GENIA Weibo MSRA Onto4 Resume Youku Ecom Mean
[…]
@@
[…]
Table 3: Comparison of prediction quality between PaDeLLM-NER and baseline methods.
Latency evaluation
@@
Prediction quality evaluation
Table 3 shows the micro F-score of PaDeLLM-NER and other baseline methods. It’s noteworthy that the micro F-scores for PaDeLLMMulti and PaDeLLMBatch are identical. Overall, PaDeLLM-NER outperforms baselines with the highest mean F-score of $84.79$ and emerging as the best-performing method in 4 out of the 9 datasets. It excels particularly in the Weibo, Youku and ACE2005 datasets and is competitive in others. We believe that the improved prediction could be attributed to shorter sequences, which lessen the challenge of long-range dependencies and enhance prediction quality. More analysis is discussed in Appendix E.
@@
In summary, the results presented in Tables 2 and 3, demonstrate that our approach not only maintains superior prediction quality but also significantly reduces inference latency.
Comparison to state-of-the-art methods
In Tables 4 and 5, we compare the micro F-scores of PaDeLLM-NER with other recent state-of-the-art methods. Our findings indicate that PaDeLLM-NER performs comparably to these methods across most datasets except for Weibo. Notably, in the Youku dataset, PaDeLLM-NER outperforms the previously best-performing method by an improvement of 1.81%. Note that the primary focus of our work is on accelerating prediction latency using LLMs compared to baseline autoregressive methods, rather than achieving state-of-the-art results.
State-of-the-art Method CoNLL03 ACE05 GENIA
BINDER Zhang et al. (2022a) 93.33 89.50 80.50
Gollie Sainz et al. (2023) 93.10 89.60 -
DeepStruct Wang et al. (2022a) 93.00 86.90 80.80
Our Method
PaDeLLM-NER 92.52 85.02 77.66
@@
Table 4: Comparing PaDeLLM-NER Performance with recent state-of-the-art methods on English datasets.
State-of-the-art Method Weibo MSRA Onto4 Resume Youku Ecom
NEZHA-BC Zhang et al. (2022b) - - - - - 82.98
SSCNN Zhang and Lu (2023) 71.81 - 82.99 96.40 86.10 81.80
W2NER Li et al. (2022) 72.32 96.10 83.08 96.65 - -
Our Method
PaDeLLM-NER 67.36 95.03* 80.81 94.98 87.91 81.85
@@
Table 5: Comparing PaDeLLM-NER Performance with recent state-of-the-art methods on Chinese datasets. “*” indicates that results are not directly comparable.
4.3 Ablation study
@@
As discussed in Section 3.1, during training, it is permissible to overlook the loss of text span “¡mention n¿”, as the model does not need to generate this specific text, which is appended during inference. However, as shown in Table 6 illustrate, omitting these texts has minimal impact on prediction quality.
One possible explanation is that during training, the more significant challenge for LLMs lies in predicting the appropriate mention texts, rather than their format. As the model can readily learns to correctly position the format “¡mention n¿”, this aspect contributes minimally to the loss computation in training. In this case, computing the loss for all text is almost equivalent to ’neglecting’ the computation of loss for “¡mention n¿”.
@@
5 Speedup Analysis
One concern noted is that batch inference does not speed up as much as inference distributed across multiple GPUs. This observation is consistent with our expectations and supported by Chen et al. (2023c) who found that batch inference in LLMs tends to be slower than single sequence inference under identical conditions, likely due to limitations in GPU memory bandwidth Cai et al. (2024).
Transitioning from these performance considerations, it’s noteworthy that PaDeLLM-NER is self-contained and can be seamlessly integrated with various generative architectures, including well-established decoder-only models Raffel et al. (2020a); Muennighoff et al. (2022); Touvron et al. (2023a, b); Bai et al. (2023); Yang et al. (2023a) and recent innovations like RWKV Peng et al. (2023), as well as multi-modal LLMs Feng et al. (2023); Liu et al. (2023b, a) for tasks like Key Information Extraction tasks Huang et al. (2019), all without needing architectural changes or additional data/modules. Also, it could be incorporated with off-the-shelf LLMs such as ChatGPT Achiam et al. (2023) and Claude-2***https://www.anthropic.com/news/claude-2 through prompt engineering without the need for further training, an aspect we plan to explore in future research.
6 Data Contamination Concerns
Since we are using LLMs as our foundational models, trained on extensive datasets from various online sources Touvron et al. (2023b); Yang et al. (2023a), there is a chance that the models may have encountered parts of our evaluation sets during their pre-training phase, albeit unintentionally. This could potentially affect our experimental results. However, the primary focus of our experiments is the comparison of our proposed method with baseline methods. Given that these methods employ the same LLM as the base model, data contamination is unlikely to significantly impact the results.
7 Conclusion
In this work, we introduce Parallel Decoding in LLM for NER (PaDeLLM-NER), a parallel decoding framework in LLMs for efficient NER. To achieve this, we recast autoregressive prediction of all label-mention pairs of traditional NER tasks into a two-step prediction: (1) predicting the number of mentions for a specific label, and (2) identifying the $n^{th}$ mention for that label. This recast allows the model to parallel decode all label-mention pairs in batches. Extensive experimental results show that the proposed method can dramatically reduces inference time, achieving inference speedup ranging from 1.76 to 10.22 times, without compromising prediction quality. Lastly, extensive ablation studies are performed to clarify the design choices of PaDeLLM-NER.
8 Limitations
Additionally, accurately counting the number of mentions remains a challenge for LLMs as discussed in Appendix E. This issue could be alleviated by implementing a specialized counting model dedicated to this task Liu and Low (2023).
Finally, there are several instances of re-computation within the pipeline that can be optimized. Specifically, input texts are encoded multiple times throughout the process. During batch decoding, certain sequences may encounter the “¡eos¿” token earlier, but due to the nature of batch inference, these sequences continue to predict. We plan to improve this in the future by implementing enhancements like KV cache reuse and batch inference with an early quit mechanism, among other strategies.
Appendix A Dataset Statistics
@@
For the MSRA dataset, we excluded four outlier instances from the test set due to their excessively high number of mentions, significantly deviating from typical examples. These outliers not only posed challenges for model inference but also risked distorting the evaluation metrics, potentially leading to an inaccurate assessment of the model’s performance on representative data.
Also, we perform label mapping to convert ground truth from special tokens to Chinese words following Lu et al. (2023a). Further details are provided in Table 10.
@@
Weibo 8
{“PER.NAM(Specific Name)”:“名称特指”, “PER.NOM(Generic Name)”:“名称代称”, “GPE.NAM(Specific Geo-Political Entity)”:“行政区特指”, “GPE.NOM(Generic Geo-Political Entity)”:“行政区代称”, “LOC.NAM(Specific Location)”:“地点特指”, “LOC.NOM(Generic Location)”:“地点代称”, “ORG.NAM(Specific Organization)”:“组织特指”, “ORG.NOM(Generic Organization)”:“组织代称” }
MSRA 3
{“LOC”:“地点’, “PER”:“名称”, “ORG”:“组织”}
OntoNotes 4.0 4
{“GPE”:“地缘”, “LOC”:“地点”, “PER”:“名称”, “ORG”:“组织”}
Resume 8
{“NAME”:“名称”, “CONT(Nationality)”:“国籍”, “RACE”:“民族”, “TITLE”:“职位”, “EDU”:“学历”, “ORG”:“公司”, “PRO(Profession)”:“专业”, “LOC(Place of Birth)”:“籍贯”}
Youku 3
{“TELEVISION”:“电视剧”, “PER(Celebrity)”:“明星”, “MISC”:“其他”}
Ecommerce 2
{“HP(brand)”:“品牌”, “HC(commodity)”:“商品”}
@@
Appendix B Reformulation Examples
@@
English text: But Fischler agreed to review his proposal after the EU ’s standing veterinary committee , mational animal health officials , questioned if such action was justified as there was only a slight risk to human health . entity type: PER ¡num¿ 1 ¡mention 1¿Fischler
Chinese 文本(text): 公报最后说,墨西哥政府认为,贩毒以及洗钱等与毒品有关的活动是威胁到国家主权和安全的一个全球性问题。(The communique concluded by stating that the Mexican government considers drug trafficking and related activities such as money laundering to be a global issue that threatens national sovereignty and security.) 指定NER标签(entity type): 地点(LOC) ¡数量¿(¡num¿) 1 ¡第1文段¿(¡mention 1¿) 墨西哥(Mexican)
@@
Appendix C Implementation Details
We train our model on all datasets for $4$ epochs, using a batch size of $128$ and a learning rate of $51e-$ 5, with the AdamW optimizer Loshchilov and Hutter (2018) and a cosine scheduler Loshchilov and Hutter (2017). The maximum input and output sequence lengths are set to 2048 and 512, respectively. Training is conducted on 8 NVIDIA A100 GPUs. This configuration is applied across all PaDeLLM-NER models, as well as two baseline models: AutoRegAug and AutoRegStruct.
Appendix D Sequence Length Reduction

### v5

@@ 4.2 Main Results
AutoReg CoNLL03 ACE05 GENIA Weibo MSRA Onto4 Resume Youku Ecom Avg.
@@
Model AI Literature Music Politics Science Movie Restaurant Avg.
SOTA
GoLLIE-7B 59.1 62.7 67.8 57.2 67.0 63.0 43.4 60.02
UniNER-7B 53.5 59.7 65.0 60.8 61.1 42.4 31.7 53.45
GLiNER-L 57.2 64.4 69.6 72.6 62.6 57.2 42.9 60.92
GNER-LLaMA-7B 63.1 68.2 75.7 69.4 69.9 68.6 47.5 66.05
@@
PaDeLLM-NER-7B 60.7 66.1 67.6 68.1 64.4 61.3 43.6 61.68
Table 3: Comparison of prediction quality with recent SOTA models in zero-shot setting.
Evaluation on inference latency
@@
Evaluation on zero-shot prediction quality
Table 3 compares the prediction quality of different models across various domains like AI, Literature, Music, Politics, Science, Movie, and Restaurant in a zero-shot setting. Among all these methods, GoLLIE-7B scores range from 43.4 in Restaurants to 67.8 in Music, with an average of 60.02. UniNER-7B has lower scores, particularly in Restaurants (31.7), and averages 53.45. GLiNER-L shows a fairly balanced performance with a high of 72.6 in Politics and an average of 60.92. GNER-LLaMA-7B excels in Music with a 75.7 score and has the highest average of all at 66.05. Our model, PaDeLLM-NER, which consistently performs well across all domains. It has the second-best average score of 61.68, following GNER-LLaMA-7B. This highlights that while it is not the top performer, it offers robust and balanced prediction capabilities across a diverse set of topics in zero-shot setting. Note that the training of PaDeLLM-NER does not incorporate the additional task scheme prompt for describing unseen entities as used in GNER-LLaMA-7B [44], which may account for the observed differences in performance.
Evaluation on supervised prediction quality
Table 4 and Table 5 present the micro F-scores of PaDeLLM-NER in comparison to other SOTA methods on supervised datasets. Notably, the micro F-scores for both PaDeLLMMulti and PaDeLLMBatch are identical. Initially, it is evident that encoder-based methods surpass LLM-based approaches, such as AutoReg and PaDeLLM-NER, within the supervised context. Nonetheless, the strength of LLM-based methods lies not in their performance under task-specific supervised settings, but rather in their superior zero-shot capabilities, which compensates for their relative shortcomings in supervised scenarios. Nevertheless, PaDeLLM-NER demonstrates SOTA performance on certain task-specific datasets, exemplified by its exceptional results on the Youku dataset.
Upon comparing PaDeLLM-NER with AutoReg, both of which are LLM-based methods, it becomes evident that PaDeLLM-NER outperforms AutoReg across both English and Chinese supervised datasets, as evidenced by its superior mean F-score. This outcome indicates that PaDeLLM-NER not only achieves lower inference latency but also maintains a higher level of prediction quality when contrasted with baseline methods.
@@
In summary, the results presented in Table 2,  34 and  5, demonstrate that our approach not only maintains superior prediction quality in both zero-shot and supervised environments but also significantly reduces inference latency.
SOTA CoNLL03 ACE05 GENIA Avg.
BINDER [50] 93.33 89.50 80.50 87.77
Gollie [30] 93.10 89.60 - -
DeepStruct [49] 93.00 86.90 80.80 86.90
AutoReg
AutoRegAug 93.08 83.04 70.16 82.09
AutoRegStruct 91.87 82.99 77.90 84.25
Ours
PaDeLLM-NER 92.52 85.02 77.66 85.07
@@
Table 4: Comparison of prediction quality with recent SOTA methods on English supervised datasets.
SOTA Weibo MSRA Onto4 Resume Youku Ecom Avg.
NEZHA-BC [54] - - - - - 82.98 -
SSCNN [55] 71.81 - 82.99 96.40 86.10 81.80 -
W2NER [63] 72.32 96.10 83.08 96.65 - - -
AutoReg
AutoRegAug 59.04 95.56* 79.20 95.80 86.07 76.02 81.94
AutoRegStruct 56.07 90.92* 80.97 95.74 86.85 81.57 82.02
Ours
PaDeLLM-NER 67.36 95.03* 80.81 94.98 87.91 81.85 84.66
@@
Table 5: Comparison of prediction quality with recent SOTA methods on English supervised datasets. “*” indicates that results are not directly comparable.
5 Speedup Analysis
One concern noted is that batch inference does not speed up as much as inference distributed across multiple GPUs. This observation is consistent with our expectations and supported by Chen et al. [19] who found that batch inference in LLMs tends to be slower than single sequence inference under identical conditions, likely due to limitations in GPU memory bandwidth [64].
Transitioning from these performance considerations, it’s noteworthy that PaDeLLM-NER is self-contained and can be seamlessly integrated with various generative architectures, including well-established decoder-only models [8, 9, 10, 11, 12, 13] and recent innovations like RWKV [65], as well as multi-modal LLMs [66, 67] for tasks like Key Information Extraction tasks [68], all without needing architectural changes or additional data/modules. Also, it could be incorporated with off-the-shelf LLMs such as ChatGPT [27] and Claude-2https://www.anthropic.com/news/claude-2 through prompt engineering without the need for further training, an aspect we plan to explore in future research.
6 Data Contamination Concerns
Since we are using LLMs as our foundational models, trained on extensive datasets from various online sources [11, 13], there is a chance that the models may have encountered parts of our evaluation sets during their pre-training phase, albeit unintentionally. This could potentially affect our experimental results. However, the primary focus of our experiments is the comparison of our proposed method with baseline methods. Given that these methods employ the same LLM as the base model, data contamination is unlikely to significantly impact the results.
7 Limitations
Additionally, accurately counting the number of mentions remains a challenge for LLMs as discussed in Appendix F. This issue could be alleviated by implementing a specialized counting model dedicated to this task [69]. Another drawback is that reformulating label-mention pairs loses location information, which hinders tasks like downstream editing. We will address this in future work. Additionally, the de-duplication mechanism is overly aggressive, potentially removing mentions that can appear under different labels—a common issue in real-world applications (see Appendix A for more details).
Finally, there are several instances of re-computation within the pipeline that can be optimized. Specifically, input texts are encoded multiple times throughout the process. During batch decoding, certain sequences may encounter the “<eos>” token earlier, but due to the nature of batch inference, these sequences continue to predict. We plan to improve this in the future by implementing enhancements like KV cache reuse and batch inference with an early quit mechanism, among other strategies.
8 Conclusion
In this study, we present PaDeLLM-NER, a parallel decoding framework for NER within LLMs. This approach enables batch parallel decoding of all label-mention pairs, significantly cutting down inference time by 1.76 to 10.22 times without sacrificing prediction accuracy.
Appendix A Ablation study
@@
As discussed in Section 3.1, during training, it is permissible to overlook the loss of text span “<mention n>”, as the model does not need to generate this specific text, which is appended during inference. However, as shown in Table 6 illustrate, omitting these texts has minimal impact on prediction quality.
One possible explanation is that during training, the more significant challenge for LLMs lies in predicting the appropriate mention texts, rather than their format. As the model can readily learns to correctly position the format “<mention n>”, this aspect contributes minimally to the loss computation in training. In this case, computing the loss for all text is almost equivalent to “neglecting” the computation of loss for “<mention n>”.
@@
We also report statistics in Table 7 and  8 showing that mentions under multiple labels are rare for both ground truth and PaDeLLM predictions. However, we recognize that the de-duplication mechanism can be overly aggressive, potentially removing mentions that appear under multiple labels—a common scenario in real-world applications. In such cases, opting not to use the de-duplication mechanism may be preferable.
Dataset Count Ratio
ACE05 1 0.00034
ConLL03 1 0.00017
GENIA 0 0
Ecom 0 0
MSRA 1 0.00013
Weibo 0 0
Youku 2 0.0012
Resume 0 0
Table 7: Mentions appear under multiple labels in ground truth.
Dataset Count Ratio
ACE05 22 0.0074
ConLL03 10 0.0017
GENIA 18 0.0034
Ecom 2 0.0012
MSRA 5 0.00089
Weibo 0 0
Youku 3 0.0019
Resume 0 0
Table 8: Mentions appear under multiple labels in PaDeLLM prediction.
Preliminary experiments for justifying the importance of two-step prediction
We conducted preliminary experiment using one-step prediction, where all mentions of the same label are predicted in a single sequence, which is referred to as OneStep in this paper. An example of OneStep parallel decoding is shown in Table 9. Note that the order of mentions is preserved as in the ground truth, following the data from the corresponding dataset. The overall latency for each example is determined by the latency of the slowest sequence. The preliminary experiment is conducted on three English dataset, i.e., CoNLL03, ACE05 and GENIA.
Entity Text NER Result
ORG <entity>ORG<text>2004-12-20T15:37:00 Microscopic microcap Everlast , mainly a maker of boxing equipment , has soared over the last several days thanks to a licensing deal with Jacques Moret allowing Moret to buy out their women ’s apparel license for \$ 30 million , on top of a \$ 12.5 million payment now . ["Microscopic microcap Everlast", "a maker of boxing equipment", "their"]
PER <entity>PER<text>2004-12-20T15:37:00 … million payment now. ["Jacques Moret", "Moret", "their", "their women"]
GPE <entity>GPE<text>2004-12-20T15:37:00 … million payment now. []
LOC <entity>LOC<text>2004-12-20T15:37:00 … million payment now. []
Table 9: Illustration of one-step parallel decoding NER approach.
Method ACE05 CoNLL03 GENIA Mean
PaDeLLMMulti 255.53 229.74 316.90 267.39
OneStepMulti 386.93 272.22 513.63 390.93
AutoRegAug 944.90 992.70 1,515.35 1150.98
AutoRegStruct 1,293.87 753.36 1,266.31 1104.51
Table 10: Comparison of inference latency.
Method ACE05 CoNLL03 GENIA Mean
PaDeLLMMulti 85.02 92.52 77.66 85.06
OneStepMulti 80.98 91.36 76.27 82.87
AutoRegAug 83.04 93.08 70.16 82.09
AutoRegStruct 82.99 91.87 77.90 84.25
Table 11: Comparison of prediction quality.
The results are reported in Table 10 and Table 11. As expected, the inference speed of one-step approach falls between that of the two-step prediction (i.e., PaDeLLM) and the purely autoregressive model. However, the prediction quality is lower compared to the two-step prediction. In other words, PaDeLLM outperforms the one-step approach in both inference speed and prediction quality, which again verifies the efficacy of PaDeLLM.
Preliminary experiments for zero-shot autoregressive baseline
We did not report zero-shot results for AutoRegaug and AutoRegstruct as they are unsuitable for this setting. Preliminary experiments show higher latency and lower F-scores compared to PaDeLLM (see Table 12 for details).
AI Literature Music Politics Science Avg
Latency (ms)
PaDeLLM 398.37 357.45 352.85 366.76 375.02 370.09
Auto_Aug 1529.95 2096.08 2545.20 2364.87 2334.05 2174.03
F-score
PaDeLLM 60.7 66.1 67.6 68.1 64.4 65.38
AutoRegAug 0.19 0.15 0.94 0.13 0.21 0.324
Table 12: Comparison of Latency and F-score between PaDeLLM and AutoRegAug under zero-shot scenarios.
Appendix B Dataset Statistics
Variant CoNLL03 ACE05 GENIA Mean
PaDeLLM-NER 92.52 85.02 77.66 85.06
+ Model scale up to 13B 93.02 84.37 78.84 85.45
Table 13: Ablations on model scaling up.
Figure 4: Percentage of different error types.
English Dataset Chinese Dataset
AutoReg CoNLL03 ACE05 GENIA Weibo MSRA Onto4 Resume Youku Ecom Mean
AutoRegAug 33.85 37.10 60.50 45.02 27.42 35.90 30.39 18.21 31.50 35.54
AutoRegStruct 28.36 49.95 49.03 62.45 18.97 25.53 53.02 18.56 22.51 36.48
Ours
PaDeLLM-NER 6.54 8.29 10.05 2.19 2.23 2.68 4.87 3.66 3.27 4.86
Table 14: Comparison of the number of generated tokens per sequence by PaDeLLM-NER with baseline methods.
@@ Appendix B Dataset Statistics
For the MSRA dataset, we excluded four outlier instances from the test set due to their excessively high number of names, significantly deviating from typical examples. These outliers not only posed challenges for model inference but also risked distorting the evaluation metrics, potentially leading to an inaccurate assessment of the model’s performance on representative data.
Also, we perform label mapping to convert ground truth from special tokens to Chinese words following [6]. Further details are provided in Table 16.
@@
Weibo 8 {“PER.NAM(Specific Name)”:“名称特指”, “PER.NOM(Generic Name)”:“名称代称”, “GPE.NAM(Specific Geo-Political Entity)”:“行政区特指”, “GPE.NOM(Generic Geo-Political Entity)”:“行政区代称”, “LOC.NAM(Specific Location)”:“地点特指”, “LOC.NOM(Generic Location)”:“地点代称”, “ORG.NAM(Specific Organization)”:“组织特指”, “ORG.NOM(Generic Organization)”:“组织代称” }
MSRA 3 {“LOC”:“地点’, “PER”:“名称”, “ORG”:“组织”}
OntoNotes 4.0 4 {“GPE”:“地缘”, “LOC”:“地点”, “PER”:“名称”, “ORG”:“组织”}
Resume 8 {“NAME”:“名称”, “CONT(Nationality)”:“国籍”, “RACE”:“民族”, “TITLE”:“职位”, “EDU”:“学历”, “ORG”:“公司”, “PRO(Profession)”:“专业”, “LOC(Place of Birth)”:“籍贯”}
Youku 3 {“TELEVISION”:“电视剧”, “PER(Celebrity)”:“明星”, “MISC”:“其他”}
Ecommerce 2 {“HP(brand)”:“品牌”, “HC(commodity)”:“商品”}
@@
Appendix C Reformulation Examples
@@
English text: But Fischler agreed to review his proposal after the EU ’s standing veterinary committee , mational animal health officials , questioned if such action was justified as there was only a slight risk to human health . entity type: PER <num> 1 <mention 1>Fischler
Chinese 文本(text): 公报最后说,墨西哥政府认为,贩毒以及洗钱等与毒品有关的活动是威胁到国家主权和安全的一个全球性问题。(The communique concluded by stating that the Mexican government considers drug trafficking and related activities such as money laundering to be a global issue that threatens national sovereignty and security.) 指定NER标签(entity type): 地点(LOC) <数量>(<num>) 1 <第1文段>(<mention 1>) 墨西哥(Mexican)
@@
Appendix D Implementation Details
We train our model on all datasets for $4$ epochs, using a batch size of $128$ and a learning rate of $51e-$ 5, with the AdamW optimizer [70] and a cosine scheduler [71]. The maximum input and output sequence lengths are set to 2048 and 512, respectively. Training is conducted on 8 NVIDIA A100 GPUs. This configuration is applied across all PaDeLLM-NER models, as well as three baseline models: AutoRegAug, AutoRegStruct as well as Onestep baseline reported in preliminary experiment. We also report the model size of each NER method in Table 18
English Method Base Language Model Chinese Method Base Language Model
BINDER BERT-base 110M NEZHA-BC NEZHA-base 110M
Gollie Code-llama 34B SSCNN not report
DeepStruct GLM10B W2NER Transformer-based 110M
AutoRegAug LLaMA-2-7B AutoRegAug Baichuan2-7B
AutoRegStruct LLaMA-2-7B AutoRegStruct Baichuan2-7B
PaDeLLM-NER LLaMA-2-7B PaDeLLM-NER Baichuan2-7B
Table 18: Model size of each NER method.
Appendix E Sequence Length Reduction